# V2A-Mapper: A Lightweight Solution for Vision-to-Audio Generation by Connecting Foundation Models

### 10 Dec 2023 v3

@@ Abstract
Building artificial intelligence (AI) systems on top of a set of foundation models (FMs) is becoming a new paradigm in AI research. Their representative and generative abilities learnt from vast amounts of data can be easily adapted and transferred to a wide range of downstream tasks without extra training from scratch. However, leveraging FMs in cross-modal generation remains under-researched when audio modality is involved. On the other hand, automatically generating semantically-relevant sound from visual input is an important problem in cross-modal generation studies. To solve this vision-to-audio (V2A) generation problem, existing methods tend to design and build complex systems from scratch using modestly sized datasets. In this paper, we propose a lightweight solution to this problem by leveraging foundation models, specifically CLIP, CLAP, and AudioLDM. We first investigate the domain gap between the latent space of the visual CLIP and the auditory CLAP models. Then we propose a simple yet effective mapper mechanism (V2A-Mapper) to bridge the domain gap by translating the visual input between CLIP and CLAP spaces. Conditioned on the translated CLAP embedding, pretrained audio generative FM AudioLDM is adopted to produce high-fidelity and visually-aligned sound. Compared to previous approaches, our method only requires a quick training of the V2A-Mapper. We further analyze and conduct extensive experiments on the choice of the V2A-Mapper and show that a generative mapper is better at fidelity and variability (FD) while a regression mapper is slightly better at relevance (CS). Both objective and subjective evaluation on two V2A datasets demonstrate the superiority of our proposed method compared to current state-of-the-art approaches - trained with 86% fewer parameters but achieving 53% and 19% improvement in FD and CS, respectively. Supplementary materials such as audio samples are provided in our demo website: https://v2a-mapper.github.io/.
@@ Introduction
Foundation models (FMs), trained on large-scale data and often making use of self-supervised learning, offer problem-agnostic representative or generative capabilities to downstream tasks via adaptation (Bommasani et al. 2021). They have demonstrated robust generalization and knowledge transfer ability across a broad spectrum of tasks in recent AI research (Zhou et al. 2023; Cao et al. 2023; Yin et al. 2023). Despite success in many uni-modal tasks spanning language (Paaß and Giesselbach 2023), vision (Awais et al. 2023), and audio (Li et al. 2023), the adaptation of FMs in problems involving multiple modalities such as cross-modal generation is greatly dominated by vision-language research (Du et al. 2022). Although attempts (Ao et al. 2022; Kreuk et al. 2023; Yang et al. 2023; Huang et al. 2023; Liu et al. 2023; Yuan et al. 2023; Ghosal et al. 2023) have been made lately to bring FMs into text-to-audio generation and achieved remarkable performance, the viability of adopting FMs in vision-to-audio generation is still unclear.
@@
In this work, we explore the feasibility of adopting foundation models in open-domain vision-to-audio generation task. As shown in Fig. 1(b), our lightweight method only requires the training of a V2A-Mapper to bridge the domain gap between the vision representative FM CLIP (Radford et al. 2021) and the audio generative FM AudioLDM (Liu et al. 2023). The V2A-Mapper is supervised by the audio representative FM CLAP (Wu et al. 2023) to learn the translation from visual space to auditory space. Leveraging the generalization and knowledge transfer ability of foundation models, the V2A-Mapper is trained with the same modestly sized dataset but the overall system can achieve much better performance. Our contribution includes: 1) investigating the potential of bringing FMs into the field of vision-to-audio generation; 2) proposing a simple but effective V2A-Mapper to connect visual and auditory FMs; 3) investigating both generative and regression strategies of the V2A-Mapper; 4) both subjective and objective evaluation on two V2A datasets demonstrate the efficiency and effectiveness of our method - it is trained with 86% fewer parameters but can achieve up to 53% and 19% improvement in fidelity (FD) and relevance (CS).
@@
AudioLDM (Liu et al. 2023) is a continuous latent diffusion model (LDM) trained in a self-supervised way with 3.3M 10-second audio clips. Conditioned on CLAP audio embedding, it generates the latent code of audio mel-spectrogram which can be decoded and converted into audio waveform. The original work only explores the usage of the LDM part in text-to-audio (T2A) generation task. Since CLAP represents text and audio jointly, AudioLDM can directly take text as input when being adapted to T2A task. We note that despite being proposed specifically for T2A generation task, AudioLDM is expected to adapt more naturally to audio features. This inspires us to ponder if we could translate a vision feature into its corresponding audio embedding in CLAP space, then we could keep AudioLDM completely intact and utilize it as an off-the-shelf audio generator FM.
@@
Table 1: Objective comparison with SOTA methods on VGGSound (video-to-sound generation) and ImageHear (image-to-sound generation). The inference time is measured as the average time spent for 100 samples through the whole pipeline from input visual prompts to output waveforms on one NVIDIA RTX A6000 GPU. Our method achieves the best on all the objective metrics.
@@
We measure the performance on two aspects, fidelity and the relevance to the visual prompt. Specifically, we use Fréchet Distance (FD) to measure the overall quality and variability of generated audio clips. FD computes the distance of embedding distributions between the synthesized and the real samples. To compare with previous methods, we also compute the Fréchet Audio Distance (FAD) (Kilgour et al. 2019). FD and FAD differ at the embedding extractor - FD uses PANNs (Kong et al. 2020) while FAD adopts VGGish (Hershey et al. 2017). Similar to  (Liu et al. 2023), we choose FD as our main evaluation metric regarding the sound quality since PANNs is superior to VGGish by considering long distance temporal change. For the relevance evaluation, we use CLIP-Score (CS) (Sheffer and Adi 2023) to get the cosine similarity between the CLIP embedding of the visual input and the Wav2CLIP (Wu et al. 2022) embedding of the generated sound. As Wav2CLIP learns an audio encoder via contrastive loss on VGGSound with the guidance of frozen CLIP image encoder, if the generated sound matches the visual input, the Wav2CLIP embedding is expected to be similar to its paired CLIP embedding.
@@
To complement the objective metrics, we also conduct a listening test to measure the fidelity of the generated audio clips and their relevance to visual prompts from a human perception perspective. We ask 20 listeners to rate audio clips of 20 randomly selected visual samples on a discrete 5-point scale in terms of fidelity and relevance, respectively. The average rating across all listeners for each algorithm is computed as Mean Opinion Score (MOS) (International Telecommunication Union 1996). We also re-code the responses into paired comparisons and infer the relative standings via indirect scaling (Agresti 1992). We calculate the degree by which other approaches exceed our method as the Just Meaningful Difference (JMD) score (e.g. a negative value indicates inferiority of other algorithms compared to ours), as elaborated more in the supplementary. More details of our human evaluation are provided in the supplementary.
@@
We use “ViT-B/32” version for CLIP model [^1] https://github.com/openai/CLIP. For CLAP model and audio generator, we use pretrained models from AudioLDM [^2] https://github.com/haoheliu/AudioLDM. For the diffusion-based V2A-Mapper, we use a cosine noise schedule with 1000 diffusion steps during training and 200 steps at inference time. We use AdamW with a learning rate of 1.1e-4, a batch size of 448 visual-audio embedding pairs, and a dropout rate of 0.1 in classifier-free guidance. We provide more implementation details including datasets used in adopted FMs, the full experiments of architecture hyperparameter tuning, and the guidance scale tuning in the supplementary section.
@@
Since AudioLDM is proposed for text-to-audio generation, the naive way of utilizing it for V2A synthesis is interleaving a captioning model to generate text input as shown in Fig. 5(a). To verify this vision-txt-audio idea, we adopt SOTA captioner BLIP (Li et al. 2022) to generate descriptions for images. For video-to-audio generation, we use the tag information provided in VGGSound. As reported in Tab. 3, although using text as bridge could mitigate the gap to some extent, it is still inferior to our method with the V2A-Mapper in both fidelity and relevance. This result indicates that the captioner is actually a bottleneck whose performance would directly affect what audio is to be generated from AudioLDM. Instead of decoding the visual condition into text format, our V2A-Mapper keeps the visual information as its latent code form and explicitly translates it from CLIP’s visual space to CLAP’s audio space, which could avoid information loss occurred during vision-txt-audio conversion. If the V2A-Mapper is skipped as illustrated in Fig. 5(b), the domain gap between vision and audio space prevents AudioLDM from generating high-fidelity and visually-relevant sound. The recent text-to-audio generation work Make-An-Audio (Huang et al. 2023) trained their audio generator with CLIP text embedding and adopted CLIP image embedding as input to handle vision-to-audio generation. Similar to the “w/o mapper” strategy, the domain gap between the visual condition and the target embedding space which their audio generator works on is not addressed. We refer readers to our demo website to observe the comparison with Make-An-Audio.
@@
Our method aims to leverage foundation models for efficient vision-to-audio generation. It can be used to enhance the immersion of human experience, such as video editing and foley design. Nevertheless, the application of this technology poses a risk if being maliciously misused on social platforms, potentially resulting in negative outcomes for society. Although significant strides have been made in audio deepfake detection research (Almutairi and Elgibreen 2022) to mitigate such concerns, the availability of ample datasets remains pivotal for improving detection accuracy. In light of this, we are committed to presenting our synthesised audio samples, intending to contribute to the advancement and fine-tuning of existing detection algorithms.
@@ Appendix A Implementation Details
All experiments are performed on a platform of four NVIDIA RTX A6000 GPUs. We train all the models with 100 epochs. The best model (diffusion w/ transformer) takes around 14 hours to train on four GPUs. Depending on architectures and strategies, training of other ablated models take from 40 minutes to 12 hours with one or four GPUs. We implement the network of V2A-Mapper in either Transformer or simple MLP. The transformer is implemented as multiple stacks of attention and feedforward layers. The number and dimension of attention heads is 12 and 64, respectively. The expansion rate for the feedforward network is set as 4. We ablate with different depths of the Transformer and empirically find 12 achieves the best result. For the simple MLP architecture, we implement it as a stack of linear projection followed by SiLU and LayerNorm. We ablate with different depths $D$ and expansion rate $E$ of intermediate layers. For the timestep embedding in diffusion models, we use learnable embedding followed by a linear projection layer to encode the scalar timestep information. Code will be released for benchmarking. For the training set of pretrained foundation models, CLIP is trained on 400M image-text pairs scrambled from internet (Radford et al. 2021). CLAP (Wu et al. 2023) is trained on 2.5M text-audio pairs including LAION-Audio-630K [^3] https://github.com/LAION-AI/audio-dataset/blob/main/laion-audio-630k/README.md, AudioSet (Gemmeke et al. 2017), AudioCaps (Kim et al. 2019), and Clotho (Drossos, Lipping, and Virtanen 2020). AudioLDM (Liu et al. 2023) is trained on 3.3M 10-second sound clips including AudioSet, AudioCaps, Freesound [^4] https://freesound.org/, and BBC Sound Effect [^5] https://sound-effects.bbcrewind.co.uk/ search.

### 14 Dec 2023 v4

@@ Abstract
Building artificial intelligence (AI) systems on top of a set of foundation models (FMs) is becoming a new paradigm in AI research. Their representative and generative abilities learnt from vast amounts of data can be easily adapted and transferred to a wide range of downstream tasks without extra training from scratch. However, leveraging FMs in cross-modal generation remains under-researched when audio modality is involved. On the other hand, automatically generating semantically-relevant sound from visual input is an important problem in cross-modal generation studies. To solve this vision-to-audio (V2A) generation problem, existing methods tend to design and build complex systems from scratch using modestly sized datasets. In this paper, we propose a lightweight solution to this problem by leveraging foundation models, specifically CLIP, CLAP, and AudioLDM. We first investigate the domain gap between the latent space of the visual CLIP and the auditory CLAP models. Then we propose a simple yet effective mapper mechanism (V2A-Mapper) to bridge the domain gap by translating the visual input between CLIP and CLAP spaces. Conditioned on the translated CLAP embedding, pretrained audio generative FM AudioLDM is adopted to produce high-fidelity and visually-aligned sound. Compared to previous approaches, our method only requires a quick training of the V2A-Mapper. We further analyze and conduct extensive experiments on the choice of the V2A-Mapper and show that a generative mapper is better at fidelity and variability (FD) while a regression mapper is slightly better at relevance (CS). Both objective and subjective evaluation on two V2A datasets demonstrate the superiority of our proposed method compared to current state-of-the-art approaches - trained with 86% fewer parameters but achieving 53% and 19% improvement in FD and CS, respectively. Supplementary materials such as audio samples are provided at our demo website: https://v2a-mapper.github.io/.
@@ Introduction
Foundation models (FMs), trained on large-scale data and often making use of self-supervised learning, offer problem-agnostic representative or generative capabilities to downstream tasks via adaptation (Bommasani et al. 2021). They have demonstrated robust generalization and knowledge transfer ability across a broad spectrum of tasks in recent AI research (Zhou et al. 2023; Cao et al. 2023; Yin et al. 2023). Despite success in many uni-modal tasks spanning language (Paaß and Giesselbach 2023), vision (Awais et al. 2023), and audio (Li et al. 2023), the adaptation of FMs in problems involving multiple modalities such as cross-modal generation is greatly dominated by vision-language research (Du et al. 2022). Although attempts (Ao et al. 2022; Kreuk et al. 2023; Yang et al. 2023; Huang et al. 2023; Liu et al. 2023a, b; Yuan et al. 2023; Ghosal et al. 2023) have been made lately to bring FMs into text-to-audio generation and achieved remarkable performance, the viability of adopting FMs in vision-to-audio generation is still unclear.
@@
In this work, we explore the feasibility of adopting foundation models in open-domain vision-to-audio generation task. As shown in Fig. 1(b), our lightweight method only requires the training of a V2A-Mapper to bridge the domain gap between the vision representative FM CLIP (Radford et al. 2021) and the audio generative FM AudioLDM (Liu et al. 2023a). The V2A-Mapper is supervised by the audio representative FM CLAP (Wu et al. 2023) to learn the translation from visual space to auditory space. Leveraging the generalization and knowledge transfer ability of foundation models, the V2A-Mapper is trained with the same modestly sized dataset but the overall system can achieve much better performance. Our contribution includes: 1) investigating the potential of bringing FMs into the field of vision-to-audio generation; 2) proposing a simple but effective V2A-Mapper to connect visual and auditory FMs; 3) investigating both generative and regression strategies of the V2A-Mapper; 4) both subjective and objective evaluation on two V2A datasets demonstrate the efficiency and effectiveness of our method - it is trained with 86% fewer parameters but can achieve up to 53% and 19% improvement in fidelity (FD) and relevance (CS).
@@
AudioLDM (Liu et al. 2023a) is a continuous latent diffusion model (LDM) trained in a self-supervised way with 3.3M 10-second audio clips. Conditioned on CLAP audio embedding, it generates the latent code of audio mel-spectrogram which can be decoded and converted into audio waveform. The original work only explores the usage of the LDM part in text-to-audio (T2A) generation task. Since CLAP represents text and audio jointly, AudioLDM can directly take text as input when being adapted to T2A task. We note that despite being proposed specifically for T2A generation task, AudioLDM is expected to adapt more naturally to audio features. This inspires us to ponder if we could translate a vision feature into its corresponding audio embedding in CLAP space, then we could keep AudioLDM completely intact and utilize it as an off-the-shelf audio generator FM.
@@
Table 1: Objective comparison with SOTA methods on VGGSound (video-to-sound generation) and ImageHear (image-to-sound generation). The inference time is measured as the average time spent for 100 samples through the whole pipeline from input visual prompts to output waveforms on one NVIDIA RTX A6000 GPU. Our method achieves the best on all the objective metrics. We also plot the comparison on VGGSound in the right diagram to showcase our method achieves better results on both metrics and contains fewer trainable parameters (smaller circle size).
@@
We measure the performance on two aspects, fidelity and the relevance to the visual prompt. Specifically, we use Fréchet Distance (FD) to measure the overall quality and variability of generated audio clips. FD computes the distance of embedding distributions between the synthesized and the real samples. To compare with previous methods, we also compute the Fréchet Audio Distance (FAD) (Kilgour et al. 2019). FD and FAD differ at the embedding extractor - FD uses PANNs (Kong et al. 2020) while FAD adopts VGGish (Hershey et al. 2017). Similar to  (Liu et al. 2023a), we choose FD as our main evaluation metric regarding the sound quality since PANNs is superior to VGGish by considering long distance temporal change. For the relevance evaluation, we use CLIP-Score (CS) (Sheffer and Adi 2023) to get the cosine similarity between the CLIP embedding of the visual input and the Wav2CLIP (Wu et al. 2022) embedding of the generated sound. As Wav2CLIP learns an audio encoder via contrastive loss on VGGSound with the guidance of frozen CLIP image encoder, if the generated sound matches the visual input, the Wav2CLIP embedding is expected to be similar to its paired CLIP embedding.
@@
To complement the objective metrics, we also conduct a listening test to measure the fidelity of the generated audio clips and their relevance to visual prompts from a human perception perspective. We ask 20 listeners to rate audio clips of 20 randomly selected visual samples on a discrete 5-point scale in terms of fidelity and relevance, respectively. The average rating across all listeners for each algorithm is computed as Mean Opinion Score (MOS) (International Telecommunication Union 1996). We also re-code the responses into paired comparisons and infer the relative standings via indirect scaling (Agresti 1992). We calculate the degree by which other approaches exceed our method as the Just Meaningful Difference (JMD) score (e.g. a negative value indicates inferiority of other algorithms compared to ours). More details of our human evaluation are provided in the supplementary.
@@
We use “ViT-B/32” version for CLIP model [^1] https://github.com/openai/CLIP. For CLAP model and audio generator, we use pretrained models from AudioLDM [^2] https://github.com/haoheliu/AudioLDM. For the diffusion-based V2A-Mapper, we use a cosine noise schedule with 1000 diffusion steps during training and 200 steps at inference time. We use AdamW with a learning rate of 1.1e-4, a batch size of 448 visual-audio embedding pairs, and a dropout rate of 0.1 in classifier-free guidance. We provide more implementation details including datasets used in adopted FMs, the full experiments of architecture hyperparameter tuning, and the guidance scale tuning in the supplementary.
@@
Since AudioLDM is proposed for text-to-audio generation, the naive way of utilizing it for V2A synthesis is interleaving a captioning model to generate text input as shown in Fig. 5(a). To verify this vision-txt-audio idea, we adopt SOTA captioner BLIP (Li et al. 2022) to generate descriptions for images. For video-to-audio generation, we use the tag information provided in VGGSound. As reported in Tab. 3, although using text as bridge could mitigate the gap to some extent, it is still inferior to our method with the V2A-Mapper in both fidelity and relevance. This result indicates that the captioner is actually a bottleneck whose performance would directly affect what audio is to be generated from AudioLDM. We provide two examples where the captioner fails at predicting correct object category in “Why Vision-Text-Audio is Bottlenecked by the Captioner” section of our demo website. We suggest readers to check the audio results to examine the bottleneck challenge. Instead of decoding the visual condition into text format, our V2A-Mapper keeps the visual information as its latent code form and explicitly translates it from CLIP’s visual space to CLAP’s audio space, which could avoid information loss occurred during vision-txt-audio conversion. If the V2A-Mapper is skipped as illustrated in Fig. 5(b), the domain gap between vision and audio space prevents AudioLDM from generating high-fidelity and visually-relevant sound. Audio examples showcasing the difference are presented in “Domain Gap Bridging Process” section of our demo website. The recent text-to-audio generation work Make-An-Audio (Huang et al. 2023) trained their audio generator with CLIP text embedding and adopted CLIP image embedding as input to handle vision-to-audio generation. Similar to the “w/o mapper” strategy, the domain gap between the visual condition and the target embedding space which their audio generator works on is not addressed. We refer readers to our demo website to observe the comparison with Make-An-Audio.
@@ Discussion
Our method aims to leverage foundation models for efficient vision-to-audio generation. It can be used to enhance the immersion of human experience, such as video editing and foley design. Nevertheless, the application of this technology poses a risk if being maliciously misused on social platforms, potentially resulting in negative outcomes for society. Although significant strides have been made in audio deepfake detection research to mitigate such concerns (Yamagishi et al. 2021), the availability of ample datasets remains pivotal for improving detection accuracy. In light of this, we are committed to presenting our synthesized audio samples, intending to contribute to the advancement and fine-tuning of existing detection algorithms.
@@ Appendix A Implementation Details
All experiments are performed on a platform of four NVIDIA RTX A6000 GPUs. We train all the models with 100 epochs. The best model (diffusion w/ transformer) takes around 14 hours to train on four GPUs. Depending on architectures and strategies, training of other ablated models take from 40 minutes to 12 hours with one or four GPUs. We implement the network of V2A-Mapper in either Transformer or simple MLP. The transformer is implemented as multiple stacks of attention and feedforward layers. The number and dimension of attention heads is 12 and 64, respectively. The expansion rate for the feedforward network is set as 4. We ablate with different depths of the Transformer and empirically find 12 achieves the best result. For the simple MLP architecture, we implement it as a stack of linear projection followed by SiLU and LayerNorm. We ablate with different depths $D$ and expansion rate $E$ of intermediate layers. For the timestep embedding in diffusion models, we use learnable embedding followed by a linear projection layer to encode the scalar timestep information. Code will be released for benchmarking. For the training set of pretrained foundation models, CLIP is trained on 400M image-text pairs scrambled from internet (Radford et al. 2021). CLAP (Wu et al. 2023) is trained on 2.5M text-audio pairs including LAION-Audio-630K [^3] https://github.com/LAION-AI/audio-dataset/blob/main/laion-audio-630k/README.md, AudioSet (Gemmeke et al. 2017), AudioCaps (Kim et al. 2019), and Clotho (Drossos, Lipping, and Virtanen 2020). AudioLDM (Liu et al. 2023a) is trained on 3.3M 10-second sound clips including AudioSet, AudioCaps, Freesound [^4] https://freesound.org/, and BBC Sound Effect [^5] https://sound-effects.bbcrewind.co.uk/ search.